# Bayesian Stability Selection and Inference on Inclusion Probabilities

### v1

@@ 4.2 Riboflavin Data
We apply Bayesian stability selection to riboflavin data. As discussed in Section 3, the data set comprises $p=4,088$ genes and the goal is to identify genes that influence riboflavin production.
Since Arashi et al. (2021) argues that there are some outliers in riboflavin data, we aim to embed a robust selection method within the stability selection framework to evaluate how a robust selector performs on these data, paying attention to outliers. Specifically, we apply Robust Least Angle Regression Selection (RLARS; Khan et al., 2007), a robust linear model selection technique available through the robustHD R package (Alfons, 2021). In the first attempt, non-informative priors are used for all genes and the number of stability selection iterations $B=1,000$. The predictor variables are standardised using the scale function prior to being input into RLARS. Genes with posterior means greater than $0.5$ are presented in Table 1. Based on Table 1, considering the posterior means and using a threshold of $\pi_{\text{thr}}=0.6$, only the gene YCKE_at is deemed stable.
@@
Gene Posterior Mean Lower Bound ($95\%$ CI) Upper Bound ($95\%$ CI)
@@
In the subsequent attempt, we prioritise the $41$ genes detected by Arashi et al. (2021) as effective genes to incorporate background knowledge from their research. We set $\tilde{\zeta}_{j}=50\%$ for these $41$ genes, indicating that this background knowledge have a $50\%$ influence on the final inference of these genes. In addition, we assume $\tilde{\xi}_{j}=70\%$ for all of them, which means that their perceived importance is $70\%$. The hyper-parameters $\tilde{\zeta}_{j}$ and $\tilde{\xi}_{j}$ can be fine-tuned by genetic experts for greater precision. Since $B=1,000$, we have $\alpha_{j}=700$ and $\beta_{j}=300$ for these $41$ genes, while for the remaining genes, we use non-informative priors with $\alpha_{j}=\beta_{j}=1$.
Genes with a posterior mean greater than $0.5$ are presented in Table 2. Based on Table 2, considering the $\pi_{\text{thr}}=0.6$ and posterior mean as the decision benchmark, Bayesian stability selection identified three stable genes using informed priors, which are YCKE_at, YOAB_at, and YXLD_at. The genes YOAB_at and YXLD_at were also identified by Bühlmann et al. (2014); however, the gene LYSC_at is replaced by YCKE_at. We interpret this as evidence that the latter is robust to outlier samples. Furthermore, it can be concluded that not all genes identified by Arashi et al. (2021) are stable when subject to data perturbations.
@@
Gene Posterior Mean Lower Bound ($95\%$ CI) Upper Bound ($95\%$ CI)
@@
The Bayesian credible intervals presented in Tables 1 and 2 have a clear interpretation; given the data and the prior knowledge, the inclusion probabilities of the variables fall into the intervals provided with probability of $0.95$.
@@
We apply Bayesian stability selection to the rat genome microarray data introduced in Section 3. Li et al. (2017) identified an outlier in the data set and subsequently removed it, focussing their analysis on the remaining data. However, we employ a robust selection method that accounts for the presence of the outlier. Following the approach used in the riboflavin data, we apply RLARS to identify the key probes across the stability selection iterations. We set the number of stability selection iterations $B=1,000$, and the probes are standardised using the scale function before being incorporated into RLARS. Initially, we use non-informative priors for all probes.
@@
Probe Posterior Mean Lower Bound ($95\%$ CI) Upper Bound ($95\%$ CI)
@@
The regularisation value $\lambda=0.0371$ is selected by using a 10-fold cross-validation and the 1-se rule on the entire data set $\mathcal{D}$. The probes are standardised using the scale function prior to processing by LASSO. Again, we use non-informative priors.
@@
Probe Posterior Mean Lower Bound ($95\%$ CI) Upper Bound ($95\%$ CI)
@@
As an example of incorporating prior knowledge, if practitioners have strong beliefs about the importance of probe 1374106_at, based on previous studies such as Huang et al. (2008), these beliefs can be represented as $\tilde{\zeta_{j}}=50\%$ and $\tilde{\xi_{j}}=100\%$. Under these values, the posterior mean of this probe is $0.605$, indicating that it is stable with the adoption of $\pi_{\text{thr}}=0.6$. Clearly, the probe 1390539_at requires less strong beliefs about its importance to be considered stable.
@@ 5 Discussion
As Christian Hennig noted in the discussion section of Meinshausen and Bühlmann (2010), it is important in practical applications to incorporate the costs of decisions regarding the inclusion or exclusion of variables into the selection process. A valuable strategy for implementing Bayesian stability selection involves integrating decision costs and conducting risk simulations based on the posterior distributions. By incorporating decision costs, trade-offs can be evaluated between different selection policies and the potential consequences of false positives and false negatives. This approach facilitates a more informed and customised variable selection process, allowing decision makers to simulate various risk scenarios and adjust their selection criteria accordingly.
In case some groups of functionally related variables are of interest, a prior can be imposed on the probability of including at least one member of a group. This approach facilitates the incorporation of background knowledge about a group into the analysis. The corresponding posterior distribution can be derived in a manner analogous to that used for individual variables, offering valuable insight into the selection behaviour of a group as a whole.
Thompson sampling (Thompson, 1933), a well-known Bayesian reinforcement learning algorithm, excels in minimising regret as an agent learns the reward distributions of, for example, a slot machine by sequentially pulling different arms, which is known as the multi-armed bandit problem. Each time the agent selects one or more arms, it updates its beliefs about the reward distributions based on the received rewards, modelled using the Beta distribution. By conceptualising the binary values of the rows of $M(\lambda)$ as the rewards associated with the selection of variables in each iteration, Bayesian stability selection can be seen as training a Thompson-sampling-based agent, where the agent gathers rewards from all arms (that is, variables) at each iteration. Unlike Thompson sampling, where the agent refrains from pulling all arms in every iteration due to associated costs, in the context of Bayesian stability selection, these costs are already accounted for by stability selection, allowing the agent to evaluate all variables simultaneously without such constraints.
One limitation of this paper is the assumption of independence in the selection of variables. This assumption allows us to model the selection status of each variable using an independent Bernoulli distribution. However, a potential improvement would be to account for the interdependencies in variable selection within the modelling process.
@@ 6 Conclusion
In this paper, we introduced a Bayesian framework for stability selection that integrates expert knowledge with data-driven insights to provide inference on inclusion probabilities. We specifically argued that selection frequencies should be interpreted as sampling probabilities and proposed a procedure to enhance these into inferential probabilities. The use of Bayesian credible intervals allows for the quantification of uncertainty regarding the selection status of variables. In addition, we demonstrated that selection frequencies can become erratic in the presence of correlated covariates, a common issue when working with biological data under the assumption that statistical correlations reflect molecular interactions.
The conjugacy of the proposed distributions removes the need for numerical estimation of posterior distributions, making Bayesian stability selection an efficient aggregation process following the stability selection iterations. We also proposed a two-step process for communication with experts, enabling the elucidation of prior distributions based on their knowledge while also allowing them to control the degree to which their opinions influence the final results.
@@
The authors declare that they have no conflict of interest.
Author contributions statement

### v2

@@ 4.2 Riboflavin Data
As discussed in Section 3, Riboflavin dataset comprises $p=4,088$ genes and the goal is to identify genes that influence riboflavin production.
Since Arashi et al. (2021) argued that there are some outliers in riboflavin data, we aim to embed a robust selection method within the stability selection framework to evaluate how a robust selector performs on these data, paying attention to outliers. Specifically, we apply Robust Least Angle Regression Selection (RLARS; Khan et al., 2007), a robust linear model selection technique available through the robustHD R package (Alfons, 2021). In the first attempt, non-informative priors are used for all genes and the number of stability selection iterations $B=1,000$. The predictor variables are standardized using the scale function prior to being input into RLARS. Genes with posterior means greater than $0.5$ are presented in Table 1. Based on Table 1, only the gene YCKE_at is deemed stable.
@@
Gene Posterior Mean Lower-Bound ($95\%$ CI) Upper-Bound ($95\%$ CI)
@@
In the subsequent attempt, we prioritize the $41$ genes detected by Arashi et al. (2021) as effective genes to incorporate background knowledge from their research. We set $\tilde{\zeta}_{j}=50\%$ for these $41$ genes, indicating that this background knowledge have a $50\%$ influence on the final inference of these genes. In addition, we assume $\tilde{\xi}_{j}=70\%$ for all of them, which means that their perceived relevance are $70\%$. The hyper-parameters $\tilde{\zeta}_{j}$ and $\tilde{\xi}_{j}$ should be fine-tuned by genetic experts for greater precision. Since $B=1,000$, we have $\alpha_{j}=700$ and $\beta_{j}=300$ for these $41$ genes, while for the remaining genes, we use non-informative priors with $\alpha_{j}=\beta_{j}=1$.
Genes with a posterior mean greater than $0.5$ are presented in Table 2. Based on Table 2, Bayesian stability selection identified three stable genes using informed priors, which are YCKE_at, YOAB_at, and YXLD_at. The genes YOAB_at and YXLD_at were also identified by Bühlmann et al. (2014); however, the gene LYSC_at is replaced by YCKE_at. We interpret this as evidence that the latter is robust to outlier samples. Furthermore, it can be concluded that not all genes identified by Arashi et al. (2021) are stable when subject to data perturbations.
@@
Gene Posterior Mean Lower-Bound ($95\%$ CI) Upper-Bound ($95\%$ CI)
@@
The Bayesian credible intervals presented in Tables 1 and 2 have a clear interpretation; given the data and the prior knowledge, the selection probabilities of the variables fall into the intervals provided with probability of $0.95$.
@@
We apply Bayesian stability selection to the rat genome microarray data introduced in Section 3. Li et al. (2017) identified an outlier in the dataset and subsequently removed it, focusing their analysis on the remaining data. However, we employ a robust selection method that accounts for the presence of the outlier. Following the approach used in the Riboflavin data, we apply RLARS to identify the key probes across the stability selection iterations. We set the number of stability selection iterations $B=1,000$, and the probes are standardized using the scale function before being incorporated into RLARS. Initially, we use non-informative priors for all probes.
@@
Probe Posterior Mean Lower-Bound ($95\%$ CI) Upper-Bound ($95\%$ CI)
@@
The regularization value $\lambda=0.0371$ is selected by using a 10-fold cross-validation and the 1-se rule on the entire dataset $\mathcal{D}$. The probes are standardized using the scale function prior to processing by Lasso. Again, we use non-informative priors.
@@
Probe Posterior Mean Lower-Bound ($95\%$ CI) Upper-Bound ($95\%$ CI)
@@
As an example of incorporating prior knowledge, if practitioners have strong beliefs about the importance of probe 1374106_at, based on previous studies such as Huang et al. (2008), these beliefs can be represented as $\tilde{\zeta_{j}}=50\%$ and $\tilde{\xi_{j}}=100\%$. Under these values, the posterior mean of this probe is $0.605$, indicating that it is stable with the adoption of $\pi_{\text{thr}}=0.6$. Clearly, the probe 1390539_at requires less strong beliefs about its relevance to be considered stable.
@@ 5 Discussion
As Christian Hennig noted in the discussion section of Meinshausen and Bühlmann (2010), it is important in practical applications to incorporate the costs of decisions regarding the inclusion or exclusion of variables into the selection process. A valuable strategy for implementing Bayesian stability selection involves integrating decision costs and conducting risk simulations based on the posterior distributions. By incorporating decision costs, trade-offs can be evaluated between different selection policies and the potential consequences of false positives and false negatives. This approach facilitates a more informed and customized variable selection process, allowing decision makers to simulate various risk scenarios and adjust their selection criteria accordingly.
In case some groups of functionally related variables are of interest, a prior can be imposed on the probability of including at least one member of a group. This approach facilitates the incorporation of background knowledge about a group into the analysis. The corresponding posterior distribution can be derived in a manner analogous to that used for individual variables, offering valuable insight into the selection behavior of a group as a whole.
Thompson sampling (Thompson, 1933), a well-known Bayesian reinforcement learning algorithm, excels in minimizing regret as an agent learns the reward distributions of, for example, a slot machine by sequentially pulling different arms, which is known as the multi-armed bandit problem. Each time the agent selects one or more arms, it updates its beliefs about the reward distributions based on the received rewards, modeled using the Beta distribution. By conceptualizing the binary values of the rows of $M(\lambda)$ as the rewards associated with the selection of variables in each iteration, Bayesian stability selection can be seen as training an expert-informed Thompson-sampling-based agent, where the agent gathers rewards from all arms (that is, variables) at each iteration. Unlike Thompson sampling, where the agent refrains from pulling all arms in every iteration due to associated costs, in the context of Bayesian stability selection, these costs are already accounted for by stability selection, allowing the agent to evaluate all variables simultaneously without such constraints.
One limitation of this paper is the assumption of independence in the selection of variables. This assumption allows us to model the selection status of each variable using an independent Bernoulli distribution. However, a potential improvement would be to account for the interdependencies in variable selection within the modeling process.
@@ 6 Conclusion
In this paper, we introduced a Bayesian framework for stability selection that integrates expert knowledge with data-driven insights to provide inference on selection probabilities. We specifically argued that selection frequencies should not be interpreted as selection probabilities and proposed a procedure to enhance selection frequencies into inferential selection probabilities. In this context, Bayesian credible intervals are used to quantify the uncertainty regarding the selection status of variables.
The conjugacy of the proposed distributions removes the need for numerical estimation of posterior distributions, making Bayesian stability selection an efficient aggregation process following the stability selection iterations. We also proposed a two-step process for communication with experts, enabling the elicitation of prior distributions based on their knowledge while also allowing them to control the degree to which their opinions influence the final results.
@@
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Authors contributions statement